OpenStudy (anonymous):

1. 2x-y=2 2. 5x+2y=14

OpenStudy (texaschic101):

you solving for x and y ?

OpenStudy (anonymous):

im trying to find the equation

OpenStudy (texaschic101):

I do not understand your question.....these are already equations

OpenStudy (texaschic101):

do you mean slope intercept form ?

OpenStudy (anonymous):

the task says "solve equation set both graphically and algebraically"

OpenStudy (the_fizicx99):

Systems of equation @texaschic101

OpenStudy (the_fizicx99):

Think you made a typo there.

OpenStudy (texaschic101):

2x - y = 2 ---> 2x - 2 = y now sub 2x - 2 in for y in the other equation 5x + 2y = 14 5x + 2(2x - 2) = 14 5x + 4x - 4 = 14 9x = 18 x = 2 now sub 2 in for x in either equation to find y 2x - y = 2 2(2) - y = 2 4 - y = 2 4 - 2 = y 2 = y check.. 2x - y = 2 2(2) - 2 = 2 4 - 2 = 2 2 = 2 (correct) solution is : (2,2) thats better

OpenStudy (texaschic101):

thats the algebraic solution

OpenStudy (the_fizicx99):

As for graphically, convert both of them into y = mx + b form and plot it onto a graph.